Using the separation of variable technique, solve: ∂u/∂x+ ∂u/∂y= 2(x+y)u
1
Expert's answer
2020-11-12T18:43:21-0500

By the separation of variables technique we write "u(x,y) = \chi(x)\phi(y)" therefore we can rewrite or equation as follows :

"\chi'(x)\phi(y) + \chi(x)\phi'(y) = 2(x+y)\chi(x)\phi(y)"

Now we will divide both parts by "u(x,y)=\chi(x)\phi(y)" and transfer all functions of a variable x to the left, and these of y to the right:

"\frac{\chi'(x)}{\chi(x)} + \frac{\phi'(y)}{\phi(y)} = 2(x+y)"

"\frac{\chi'(x)}{\chi(x)}-2x = 2y-\frac{\phi'(y)}{\phi(y)}"

On the left we have a function of x only and on the right of y only. Therefore both parts of this equality must be a constant (as x,y are independent) that we will call "\lambda" . Now let's solve the equations for "\phi(y), \chi(x)" :

"\frac{\chi'(x)}{\chi(x)}-2x=\lambda"

"\chi'(x) = (2x+\lambda)\chi(x)"

"\chi(x) = A_\lambda e^{x^2+\lambda x}"


"2y-\frac{\phi'(y)}{\phi(y)}=\lambda"

"\phi'(y) = (2y-\lambda)\phi(y)"

"\phi(y)=B_\lambda e^{y^2-\lambda y}"


"u_\lambda(x,y)=\phi_\lambda(y)\times\chi_\lambda(x)=C_\lambda e^{x^2+y^2+\lambda(x-y)}"

The range of possible "\lambda" , the constants "C_\lambda" are determined by boundary and initial conditions, the general sum can be expressed as a linear (possibly infinite) combination of "u_\lambda(x,y)" for permitted values of "\lambda" .